OpenStudy (anonymous):

What is a radius of a circle whose equation is x^2+y^2=9

OpenStudy (larseighner):

What is the definition of a circle?

OpenStudy (ikram002p):

the equation of center (0,0) with raduis r has the equation of :- \( x^2+y^2=r^2\)

OpenStudy (larseighner):

@ikram002p is right. Here is how you get that: A circle is defined as all the points in a plane that are the same distance from a given point in the plane. Now given the point (0,0), what is the distance r to an arbitrary point (x, y)? The difference in the first coordinate (abscissa) values is x-0. That is the length of one leg of a right triangle. The difference in the the second coordinates is the length of the other leg of a right triangle, so thanks to Pythagoras the distance between the arbitrary point (x, y) and the origin is: \[\large r = \sqrt{(x-0)^2 + (y-0)^2} = \sqrt{{x^2+y^2}}\] Squaring both sides: \[\large r^2 = x^2+y^2\]